# Signed Networks: theory, methods, and applications

### v1

@@ 3.1 Heiderian (cognitive) balance
We first explore Heider’s notion of balance, which we refer to as cognitive balance. In his foundational work [heiderAttitudesCognitiveOrganization1946], Heider analysed systems composed of two or three elements, which can be either people (denoted by $p$, $o$, $q$) or impersonal elements like objects or concepts (denoted by $x$). He distinguished between two types of pairwise relations between elements: unit relations and attitudes. A unit relation, written as $pUq$, refers to an objective connection between $p$ and $q$ that goes beyond personal perception, such as belonging, similarity, or spatial proximity. These relations are symmetric, even if not equivalent in meaning from both directions (for instance, $p$ belongs to $q$ always implies that $q$ contains $p$). In contrast, an attitude, written as $pLq$, expresses a subjective evaluation like liking or trusting. Unlike unit relations, attitudes can be asymmetric because they are rooted in the local individual perception. Importantly, any relation $U$ or $L$ can be negated, so that, if $L$ means liking, $p\neg Lq$ means that $p$ does not like $q$. In summary, the differentiation between unit relations and attitudes refers to the differentiation between the material configuration and the psychological configuration between two or more elements.
@@
Moreover, it is worth mentioning that some studies have focused on empirically validating the existence of this “subjective experience”, namely, the perception that certain configurations feel more tense or less stable, and its limitations. For instance, [morrissetteExperimentalStudyTheory1958] assesses stress levels of respondents in certain hypothetical configurations of a system, and [whiteLimitationBalanceTheory1977] shows the importance of individuals feeling personally involved in the situation to test any hypothesis related to Heider’s balance. Other examples that explore this cognitive perspective of balance in hypothetical situations are [millerStructuralBalanceDyads1972][horowitzInductionForcesDiscussion1951][chiangTriadicBalanceBrain2020], which explores with neuroimaging differences in the brain when people are in unbalanced vs. balanced situations, or [jordanBehavioralForcesThat1953], which explores the existence of conflict aversion beyond balance.
Heider’s notion of balance has some limitations [cartwrightStructuralBalanceGeneralization1956]. First, negative relations in his framework are ambiguous, as they can represent either the absence of a positive relation or an explicitly antagonistic relation. For instance, $p\neg Lq$ could mean that $p$ simply does not consider $q$ a friend, or that $p$ and $q$ are actual enemies. The notion of balance in the system as a whole is ambiguous itself: a single directional positive link can be balanced with respect to transitivity but unbalanced with respect to reciprocity. Second, Heider’s formulation is not trivially generalizable beyond pairs and triads, while most social or complex networks involve larger groups of interacting entities. Finally, his work was restricted to cognitive or attitudinal relations between elements, even though the underlying ideas had the potential to be generalised to arbitrary signed relations.
@@ 3.2 Hararian (structural) balance
Heider’s notion of balance is grounded in the individual and local perceptions of people within a system of two or three elements. In his framework, balance is described conceptually rather than through formal rules. In this setting, the key insight of Cartwright and Harary was to recognise that Heider’s theory, though expressed in psychological and local terms, pointed towards general structural properties that could be formalised within the language of graph theory [cartwrightStructuralBalanceGeneralization1956] [^8] Harary first introduced structural balance as a purely mathematical property of graphs [hararyNotionBalanceSigned1953]. It was only three years later that he and Cartwright fully developed the connections between this mathematical notion of balance and Heider’s theory [cartwrightStructuralBalanceGeneralization1956].. This new conceptualisation of balance, often referred to as Hararian or structural balance, quickly gained popularity and has, over time, somewhat overshadowed Heider’s original and broader notion of balance.
Cartwright and Harary’s core realization was that, within Heider’s framework, if (1) we restrict attention to systems in which relations represent only attitudes, ignoring unit relations, (2) disregard the directionality of attitudes so that liking or disliking is treated as mutual and, importantly, no longer local [^9] In their original work [cartwrightStructuralBalanceGeneralization1956], the authors considered directed graphs; however, most applications of structural balance theory are concerned with undirected networks., and (3) evaluate whether a triad is balanced from a global perspective (that is, disregarding individual perceptions and considering the system as a whole), then balanced triads are precisely those containing an even number of negative relations, whereas unbalanced triads contain an odd number. They then extended this principle beyond triads, proposing that the same rule could apply to any cycle in a graph: a cycle is balanced if it has an even number of negative edges, and unbalanced if the number is odd. Finally, they generalised this to networks by defining a network as balanced if and only if all its cycles are balanced. With these steps, we arrive at the following definition:
@@ 3.3 Comparison between Heiderian and Hararian balance
Because of these differences, applying one framework in contexts suited for the other can be problematic, and it can lead to contradicting predictions. For example, conflict aversion, a form of Heiderian balance, may favour dissolving negative links, even if they take part in balanced cycles in Harary’s sense. Unfortunately, empirical studies have blurred the line between the two, often interpreting results that support Heider’s cognitive mechanisms as evidence for structural balance in the Hararian sense or viceversa. This conflation is partly historical: Cartwright and Harary originally presented structural balance as a formal generalisation of Heider’s ideas, which encouraged researchers to treat them as equivalent despite their conceptual divergence. A constructive way forward is to recognise that these approaches can complement rather than compete with each other. An example is the work of Doreian and Hummon [doreianEvolutionHumanSigned2004, hummonDynamicsSocialBalance2003], who developed a dynamic model that integrates both perspectives: Heiderian balance at the local cognitive level and structural balance at the group level. Most of the focus in this review will be on Harary’s structural balance, although Heiderian balance remains important and can be useful in certain contexts.
@@ 3.4 Mathematical properties of balanced networks
Harary’s mathematical definition of balance has the advantage of unlocking the full machinery of graph theory. With this framework, we can derive results and properties that hold for any signed graph, regardless of its size or the details of the complex system it represents. One of the most striking examples of this generality is Harary’s theorem [hararyNotionBalanceSigned1953], which states that any balanced graph, regardless of its size, can be divided into two antagonistic groups.
@@
In other words, the defining property of balanced networks is that they admit a natural bipartition into two factions, often referred to as the balanced factions. Thus, in a balanced network, every node has either a relation of effective alliance or enmity with any other node in the network, with no ambiguous relations. This immediately makes tasks like link sign prediction trivial: for any pair of disconnected nodes $i,j$ in a balanced graph, any structural-balance-based algorithm will predict a positive connection when the nodes belong to the same balanced faction, and negative otherwise (see Section 4.3 for an overview on sign prediction methods). Furthermore, structural balance has strong algebraic consequences. First, the adjacency matrix $A$ and its unsigned counterpart $|A|$ are cospectral; that is, they have identical eigenvalue spectra [acharyaSpectralCriterionCycle1980]. Second, any balanced signed graph can be transformed into an equivalent unsigned graph through a node-based operation called switching [zaslavskySignedGraphs1982], which reverses the signs of all edges incident to selected nodes. This means that many structural and dynamical properties of balanced signed networks can be studied using the rich theory of unsigned graphs [zaslavskyMatricesTheorySigned2013].
@@
With respect to dynamics, balance guarantees a coherent propagation of information. Coherent means that the effect of a signal does not depend on the particular path it takes: because all paths between any two nodes have the same sign, any piece of information arrives at the target node with the same polarity, independently of the chosen path. This typically gives rise to polarised dynamical states, where the dynamical states of nodes belonging to opposite balanced factions have opposite sign [altafiniConsensusProblemsNetworks2013] (see Section 5.1).
@@
In terms of energy minimisation, balanced configurations correspond to ground states of the Ising Hamiltonian [^10] In statistical physics, the Ising model describes interacting spins that can take two values, +1+1 or 1-1. The Hamiltonian encodes the energy of a configuration, and ground states are those with minimal energy. Interpreting factions as spin values, a balanced partition minimises the Hamiltonian, so balance corresponds to an energy minimum.. This provides a solid connection between structural balance and statistical physics, and more concretely, the theory of spin glasses [facchettiComputingGlobalStructural2011, facchettiExploringLowenergyLandscape2012]. For instance, one can use simulated annealing techniques to find the balanced factions, or even extend these techniques to unbalanced networks that are close to a balanced state [traagCommunityDetectionNetworks2009]. The energy-minimisation interpretation of balance also provides a suggestive explanation as to why many social and biological systems gravitate toward balanced structures [facchettiComputingGlobalStructural2011].
@@
Harary’s characterisation of balance has an exact counterpart, known as antibalance [hararyStructuralDuality1957]. A graph is antibalanced if it becomes balanced after flipping the sign of every edge. Equivalently, an antibalanced graph satisfies that every cycle has an even number of positive edges. In matrix terms, antibalance means that the spectrum of $A$ is the negative of the spectrum of $|A|$. Antibalanced networks can also be partitioned into two antibalanced factions, so that negative edges lie within each group and positive edges across the two groups.
@@
Arguably, the most important criticism of Harary’s theory is that it is too restrictive for many practical applications. In reality, most signed networks do not fulfil the conditions for strict balance. In some cases, however, they do satisfy a weaker form of balance, whereby nodes can be divided into more than two mutually antagonistic factions. This observation motivated Davis to introduce the notion of weak balance [davisClusteringStructuralBalance1967] [^11] Davis originally termed this concept “clustering”; today “clustering” typically refers to community‑detection algorithms or clustering coefficients.: A signed graph is weakly balanced if it contains no cycle with exactly one negative edge [^12] To distinguish the two notions of structural balance, Harary’s original definition is often called strong balance.. Weakly balanced graphs fulfil an analogue of Harary’s theorem, called Davis’s structure theorem: a signed graph is weakly balanced if and only if its nodes can be partitioned into $k\geq 2$ disjoint factions such that every positive edge lies within a faction and every negative edge lies between factions (see Figure 7). In this sense, weak balance generalises Harary’s bipartition by allowing more than two balanced factions, thus capturing multipolar configurations (e.g., multiple rival coalitions) that strong balance cannot.
@@
More recently, Doreian and Mrvar [doreianPartitioningSignedSocial2009] argued that even weak balance remains too rigid for many empirical networks. Accordingly, they proposed the relaxed structural balance block model. A signed graph is relaxed-balanced if its adjacency matrix can be permuted into a block structure containing only three block types: positive (all links are positive), negative (all links are negative), or null (no links), without restriction on block location (see Figure 7). Relaxed balance accounts for phenomena invisible under strict or weak balance, such as mediation (actors bridging hostile groups), differential popularity (actors receiving uniformly positive links), and internal hostility (groups where all members dislike each other).
@@ 3.5 Balance indices
The simplest way to measure balance uses motifs. It relies on the idea that small subgraphs can serve as proxies for global structure. Harary himself pioneered this approach, proposing cycles of size order $N$ as a reasonable indicator of the level of balance [harary1955local]. Today, triangles are the most common choice, since they are the smallest cycles that can reveal balance. The idea is to compare how many of them are balanced and how many are not [galloTestingStructuralBalance2024, haoProperNetworkRandomization2023, arefMultilevelStructuralEvaluation2020]. A convenient way to do so is computing the trace of the cube of the adjacency matrix, $\mathrm{tr}(A^{3})$. To control for the total number of triangles, one can use the ratio
@@
The frustration-based approach [facchettiComputingGlobalStructural2011, facchettiExploringLowenergyLandscape2012, aref2018measuring, arefBalanceFrustrationSigned2019, arefMultilevelStructuralEvaluation2020], is rooted in spin glass theory and statistical mechanics [mezardSpinGlassTheory1987], and treats the network as a system of interacting spins. Such systems can exhibit frustration, a phenomenon that occurs when no configuration of spins can satisfy all interactions at once. In other words, no matter how the spins are aligned, at least one positive edge connects spins that point in opposite directions, or at least one negative edge connects spins that point in the same direction. Consider, for example, a triangle where all edges are negative: each pair of nodes prefers to point in opposite directions, but three spins cannot all be opposite, so at least one interaction must remain unsatisfied. Even in the most favourable configuration, understood as the one that satisfies as many interactions as possible, some edges remain in conflict (frustrated). The minimum number of such edges is called the frustration index of the network. Equivalently, the frustration index is the minimum number of edges that must be removed or have their sign flipped to obtain a system with no frustration.
@@
Frustration-based approaches have several advantages. They establish a direct link between imbalance and a physically meaningful quantity, namely, the number of unsatisfied interactions. They also allow us to use theoretical concepts from spin glass physics, such as energy landscapes and optimisation principles, as well as powerful computational tools originally developed in statistical mechanics and combinatorial optimisation. Their main limitation lies in computation. Determining the frustration index is an NP-hard problem, so exact solutions are not feasible for large networks, and one must rely on heuristic or approximate algorithms [facchettiComputingGlobalStructural2011].
The algebraic approach assesses balance through spectral properties of signed matrices. The most common choice is the opposing Laplacian (see Section 2.3). Just as the second smallest eigenvalue of the unsigned Laplacian measures connectivity in an unsigned network, the smallest eigenvalue $\mu_{1}$ of the opposing Laplacian serves as an indicator of imbalance [kunegisSpectralAnalysisSigned2010]. It is equal to zero if and only if the network is perfectly balanced, while larger values indicate an increasingly unbalanced structure. An advantage of this spectral viewpoint is that it connects structural balance to network dynamics such as synchronisation, diffusion, or consensus processes. In particular, $\mu_{1}$ controls the rate at which diffusive dynamics would converge to the trivial state where all node variables vanish [altafiniConsensusProblemsNetworks2013]. Highly unbalanced networks have large $\mu_{1}$ and therefore decay rapidly, whereas perfectly balanced networks never decay to that state at all.
The last approach we will discuss uses closed walks as proxies for cycles [^13] This assumption has proved valid for a wide range of empirical systems [talagaPolarizationMultiscaleStructural2023]. [kirkleyBalanceSignedNetworks2019, estradaWalkbasedMeasureBalance2014, estradaRethinkingStructuralBalance2019]. It compares the number of positive and negative closed walks through powers of the adjacency matrix (Section 2.1), weighting their contributions by length. This is done by assigning a penalty $c_{k}$ to walks of length $k$:
@@
A value $\kappa=1$ indicates perfect balance [^14] This follows from the fact that $A$ and $|A|$ are cospectral if and only if the network is balanced. Several choices of $c_{k}$ appear in the literature, such as $c_{k}=\alpha^{k}$ [singhMeasuringBalanceSigned2017] or $c_{k}=\alpha^{k}/k$ [kirkleyBalanceSignedNetworks2019]. The most common choice is $c_{k}=\beta^{k}/k!$, where $\beta>0$ controls the penalty for longer walks [estradaWalkbasedMeasureBalance2014]. With this choice, Eq. (2) becomes the matrix exponential $\mathrm{tr}(e^{\beta A})$.
This framework admits a node-level extension that quantifies the contribution of each node to global balance [diaz-diazMathematicalModelingLocal2024, diazdiazthesis]:
@@
The walk-based approach has notable advantages. Closed walks can be counted efficiently using matrix powers, and the matrix exponential formulation allows us to leverage well-established theoretical properties and efficient numerical algorithms [molerNineteenDubiousWays2003]. It also creates links with statistical mechanics, where $\beta$ plays the role of an inverse temperature, and with diffusion processes, where $\beta$ represents time [estrada2012Physics]. However, closed walks have an important limitation: they allow backtracking, that is, repeatedly moving back and forth along the same edge instead of exploring genuine cycles. As a result, long backtracking walks may dominate shorter ones, even when penalisation factors are applied. Recent advances have proposed principled methods to mitigate this effect with remarkable success [talagaPolarizationMultiscaleStructural2023].

### v2

@@ 3.1 Heiderian (cognitive) balance
We first explore Heider’s notion of balance, which we refer to as cognitive balance. In his foundational work [177], Heider analysed systems composed of two or three elements, which can be either people (denoted by $p$, $o$, $q$) or impersonal elements like objects or concepts (denoted by $x$). He distinguished between two types of pairwise relations between elements: unit relations and attitudes. A unit relation, written as $pUq$, refers to an objective connection between $p$ and $q$ that goes beyond personal perception, such as belonging, similarity, or spatial proximity. These relations are symmetric, even if not equivalent in meaning from both directions (for instance, $p$ belongs to $q$ always implies that $q$ contains $p$). In contrast, an attitude, written as $pLq$, expresses a subjective evaluation like liking or trusting. Unlike unit relations, attitudes can be asymmetric because they are rooted in the local individual perception. Importantly, any relation $U$ or $L$ can be negated, so that, if $L$ means liking, $p\neg Lq$ means that $p$ does not like $q$. In summary, the differentiation between unit relations and attitudes refers to the differentiation between the material configuration and the psychological configuration between two or more elements.
@@
Moreover, it is worth mentioning that some studies have focused on empirically validating the existence of this “subjective experience”, namely, the perception that certain configurations feel more tense or less stable, and its limitations. For instance, [279] assesses stress levels of respondents in certain hypothetical configurations of a system, and [397] shows the importance of individuals feeling personally involved in the situation to test any hypothesis related to Heider’s balance. Other examples that explore this cognitive perspective of balance in hypothetical situations are [272][185][70], which explores with neuroimaging differences in the brain when people are in unbalanced vs. balanced situations, or [203], which explores the existence of conflict aversion beyond balance.
Heider’s notion of balance has some limitations [55]. First, negative relations in his framework are ambiguous, as they can represent either the absence of a positive relation or an explicitly antagonistic relation. For instance, $p\neg Lq$ could mean that $p$ simply does not consider $q$ a friend, or that $p$ and $q$ are actual enemies. The notion of balance in the system as a whole is ambiguous itself: a single directional positive link can be balanced with respect to transitivity but unbalanced with respect to reciprocity. Second, Heider’s formulation is not trivially generalizable beyond pairs and triads, while most social or complex networks involve larger groups of interacting entities. Finally, his work was restricted to cognitive or attitudinal relations between elements, even though the underlying ideas had the potential to be generalised to arbitrary signed relations.
@@ 3.2 Hararian (structural) balance
Heider’s notion of balance is grounded in the individual and local perceptions of people within a system of two or three elements. In his framework, balance is described conceptually rather than through formal rules. In this setting, the key insight of Cartwright and Harary was to recognise that Heider’s theory, though expressed in psychological and local terms, pointed towards general structural properties that could be formalised within the language of graph theory [55] [^8] Harary first introduced structural balance as a purely mathematical property of graphs [168]. It was only three years later that he and Cartwright fully developed the connections between this mathematical notion of balance and Heider’s theory [55].. This new conceptualisation of balance, often referred to as Hararian or structural balance, quickly gained popularity and has, over time, somewhat overshadowed Heider’s original and broader notion of balance.
Cartwright and Harary’s core realization was that, within Heider’s framework, if (1) we restrict attention to systems in which relations represent only attitudes, ignoring unit relations, (2) disregard the directionality of attitudes so that liking or disliking is treated as mutual and, importantly, no longer local [^9] In their original work [55], the authors considered directed graphs; however, most applications of structural balance theory are concerned with undirected networks., and (3) evaluate whether a triad is balanced from a global perspective (that is, disregarding individual perceptions and considering the system as a whole), then balanced triads are precisely those containing an even number of negative relations, whereas unbalanced triads contain an odd number. They then extended this principle beyond triads, proposing that the same rule could apply to any cycle in a graph: a cycle is balanced if it has an even number of negative edges, and unbalanced if the number is odd. Finally, they generalised this to networks by defining a network as balanced if and only if all its cycles are balanced. With these steps, we arrive at the following definition:
@@ 3.3 Comparison between Heiderian and Hararian balance
Because of these differences, applying one framework in contexts suited for the other can be problematic, and it can lead to contradicting predictions. For example, conflict aversion, a form of Heiderian balance, may favour dissolving negative links, even if they take part in balanced cycles in Harary’s sense. Unfortunately, empirical studies have blurred the line between the two, often interpreting results that support Heider’s cognitive mechanisms as evidence for structural balance in the Hararian sense or viceversa. This conflation is partly historical: Cartwright and Harary originally presented structural balance as a formal generalisation of Heider’s ideas, which encouraged researchers to treat them as equivalent despite their conceptual divergence. A constructive way forward is to recognise that these approaches can complement rather than compete with each other. An example is the work of Doreian and Hummon [100, 195], who developed a dynamic model that integrates both perspectives: Heiderian balance at the local cognitive level and structural balance at the group level. Most of the focus in this review will be on Harary’s structural balance, although Heiderian balance remains important and can be useful in certain contexts.
@@ 3.4 Mathematical properties of balanced networks
Harary’s mathematical definition of balance has the advantage of unlocking the full machinery of graph theory. With this framework, we can derive results and properties that hold for any signed graph, regardless of its size or the details of the complex system it represents. One of the most striking examples of this generality is Harary’s theorem [168], which states that any balanced graph, regardless of its size, can be divided into two antagonistic groups.
@@
In other words, the defining property of balanced networks is that they admit a natural bipartition into two factions, often referred to as the balanced factions. Thus, in a balanced network, every node has either a relation of effective alliance or enmity with any other node in the network, with no ambiguous relations. This immediately makes tasks like link sign prediction trivial: for any pair of disconnected nodes $i,j$ in a balanced graph, any structural-balance-based algorithm will predict a positive connection when the nodes belong to the same balanced faction, and negative otherwise (see Section 4.3 for an overview on sign prediction methods). Furthermore, structural balance has strong algebraic consequences. First, the adjacency matrix $A$ and its unsigned counterpart $|A|$ are cospectral; that is, they have identical eigenvalue spectra [1]. Second, any balanced signed graph can be transformed into an equivalent unsigned graph through a node-based operation called switching [412], which reverses the signs of all edges incident to selected nodes. This means that many structural and dynamical properties of balanced signed networks can be studied using the rich theory of unsigned graphs [413].
@@
With respect to dynamics, balance guarantees a coherent propagation of information. Coherent means that the effect of a signal does not depend on the particular path it takes: because all paths between any two nodes have the same sign, any piece of information arrives at the target node with the same polarity, independently of the chosen path. This typically gives rise to polarised dynamical states, where the dynamical states of nodes belonging to opposite balanced factions have opposite sign [9] (see Section 5.1).
@@
In terms of energy minimisation, balanced configurations correspond to ground states of the Ising Hamiltonian [^10] In statistical physics, the Ising model describes interacting spins that can take two values, +1+1 or 1-1. The Hamiltonian encodes the energy of a configuration, and ground states are those with minimal energy. Interpreting factions as spin values, a balanced partition minimises the Hamiltonian, so balance corresponds to an energy minimum.. This provides a solid connection between structural balance and statistical physics, and more concretely, the theory of spin glasses [113, 112]. For instance, one can use simulated annealing techniques to find the balanced factions, or even extend these techniques to unbalanced networks that are close to a balanced state [386]. The energy-minimisation interpretation of balance also provides a suggestive explanation as to why many social and biological systems gravitate toward balanced structures [113].
@@
Harary’s characterisation of balance has a counterpart, known as antibalance [170]. A graph is antibalanced if it becomes balanced after flipping the sign of every edge. Equivalently, an antibalanced graph satisfies that every cycle has an even number of positive edges. In matrix terms, antibalance means that the spectrum of $A$ is the negative of the spectrum of $|A|$. Antibalanced networks can also be partitioned into two antibalanced factions, so that negative edges lie within each group and positive edges across the two groups.
@@
Arguably, the most important criticism of Harary’s theory is that it is too restrictive for many practical applications. In reality, most signed networks do not fulfil the conditions for strict balance. In some cases, however, they do satisfy a weaker form of balance, whereby nodes can be divided into more than two mutually antagonistic factions. This observation motivated Davis to introduce the notion of weak balance [79] [^11] Davis originally termed this concept “clustering”; today “clustering” typically refers to community‑detection algorithms or clustering coefficients.: A signed graph is weakly balanced if it contains no cycle with exactly one negative edge [^12] To distinguish the two notions of structural balance, Harary’s original definition is often called strong balance.. Weakly balanced graphs fulfil an analogue of Harary’s theorem, called Davis’s structure theorem: a signed graph is weakly balanced if and only if its nodes can be partitioned into $k\geq 2$ disjoint factions such that every positive edge lies within a faction and every negative edge lies between factions (see Figure 7). In this sense, weak balance generalises Harary’s bipartition by allowing more than two balanced factions, thus capturing multipolar configurations (e.g., multiple rival coalitions) that strong balance cannot.
@@
More recently, Doreian and Mrvar [97] argued that even weak balance remains too rigid for many empirical networks. Accordingly, they proposed the relaxed structural balance block model. A signed graph is relaxed-balanced if its adjacency matrix can be permuted into a block structure containing only three block types: positive (all links are positive), negative (all links are negative), or null (no links), without restriction on block location (see Figure 7). Relaxed balance accounts for phenomena invisible under strict or weak balance, such as mediation (actors bridging hostile groups), differential popularity (actors receiving uniformly positive links), and internal hostility (groups where all members dislike each other).
@@ 3.5 Balance indices
The simplest way to measure balance uses motifs. It relies on the idea that small subgraphs can serve as proxies for global structure. Harary himself pioneered this approach, proposing cycles of size order $N$ as a reasonable indicator of the level of balance [169]. Today, triangles are the most common choice, since they are the smallest cycles that can reveal balance. The idea is to compare how many of them are balanced and how many are not [138, 167, 18]. A convenient way to do so is computing the trace of the cube of the adjacency matrix, $\mathrm{tr}(A^{3})$. To control for the total number of triangles, one can use the ratio
@@
The frustration-based approach [113, 112, 21, 22, 18], is rooted in spin glass theory and statistical mechanics [269], and treats the network as a system of interacting spins. Such systems can exhibit frustration, a phenomenon that occurs when no configuration of spins can satisfy all interactions at once. In other words, no matter how the spins are aligned, at least one positive edge connects spins that point in opposite directions, or at least one negative edge connects spins that point in the same direction. Consider, for example, a triangle where all edges are negative: each pair of nodes prefers to point in opposite directions, but three spins cannot all be opposite, so at least one interaction must remain unsatisfied. Even in the most favourable configuration, understood as the one that satisfies as many interactions as possible, some edges remain in conflict (frustrated). The minimum number of such edges is called the frustration index of the network. Equivalently, the frustration index is the minimum number of edges that must be removed or have their sign flipped to obtain a system with no frustration.
@@
Frustration-based approaches have several advantages. They establish a direct link between imbalance and a physically meaningful quantity, namely, the number of unsatisfied interactions. They also allow us to use theoretical concepts from spin glass physics, such as energy landscapes and optimisation principles, as well as powerful computational tools originally developed in statistical mechanics and combinatorial optimisation. Their main limitation lies in computation. Determining the frustration index is an NP-hard problem, so exact solutions are not feasible for large networks, and one must rely on heuristic or approximate algorithms [113].
The algebraic approach assesses balance through spectral properties of signed matrices. The most common choice is the opposing Laplacian (see Section 2.3). Just as the second smallest eigenvalue of the unsigned Laplacian measures connectivity in an unsigned network, the smallest eigenvalue $\mu_{1}$ of the opposing Laplacian serves as an indicator of imbalance [224]. It is equal to zero if and only if the network is perfectly balanced, while larger values indicate an increasingly unbalanced structure. An advantage of this spectral viewpoint is that it connects structural balance to network dynamics such as synchronisation, diffusion, or consensus processes. In particular, $\mu_{1}$ controls the rate at which diffusive dynamics would converge to the trivial state where all node variables vanish [9]. Highly unbalanced networks have large $\mu_{1}$ and therefore decay rapidly, whereas perfectly balanced networks never decay to that state at all.
The last approach we will discuss uses closed walks as proxies for cycles [^13] This assumption has proved valid for a wide range of empirical systems [374]. [212, 106, 109]. It compares the number of positive and negative closed walks through powers of the adjacency matrix (Section 2.1), weighting their contributions by length. This is done by assigning a penalty $c_{k}$ to walks of length $k$:
@@
A value $\kappa=1$ indicates perfect balance [^14] This follows from the fact that $A$ and $|A|$ are cospectral if and only if the network is balanced. Several choices of $c_{k}$ appear in the literature, such as $c_{k}=\alpha^{k}$ [356] or $c_{k}=\alpha^{k}/k$ [212]. The most common choice is $c_{k}=\beta^{k}/k!$, where $\beta>0$ controls the penalty for longer walks [106]. With this choice, Eq. (2) becomes the matrix exponential $\mathrm{tr}(e^{\beta A})$.
This framework admits a node-level extension that quantifies the contribution of each node to global balance [90, 93]:
@@
The walk-based approach has notable advantages. Closed walks can be counted efficiently using matrix powers, and the matrix exponential formulation allows us to leverage well-established theoretical properties and efficient numerical algorithms [276]. It also creates links with statistical mechanics, where $\beta$ plays the role of an inverse temperature, and with diffusion processes, where $\beta$ represents time [107]. However, closed walks have an important limitation: they allow backtracking, that is, repeatedly moving back and forth along the same edge instead of exploring genuine cycles. As a result, long backtracking walks may dominate shorter ones, even when penalisation factors are applied. Recent advances have proposed principled methods to mitigate this effect with remarkable success [374].